# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502 Japan

### v2

@@ Abstract
Aharony and Fisher showed that non-local dipolar effects in magnetism destabilize the Heisenberg fixed point in real ferromagnets, leading to a new fixed point, called the dipolar fixed point. The non-perturbative nature of the new fixed point, however, has not been uncovered for many decades. Inspired by the recent understanding that the dipolar fixed point is scale-invariant but not conformal invariant, we perform the Monte Carlo simulation of the local Heisenberg-dipolar model on the lattice of $40^{3}$ by introducing the local cost function parameterized by a parameter $\lambda$, and study its critical exponents, which should become identical to the dipolar fixed point of Aharony and Fisher in the infinite coupling limit $\lambda=\infty$. We find that the critical exponents become noticeably different from those of the Heisenberg fixed point for a finite coupling constant $\lambda=8$ (e.g. $\nu=0.601(2)(^{+0}_{-2})$ in the local Heisenberg-dipolar model while $\nu=0.712(1)(^{+3}_{-0})$ in the Heisenberg model), and the spin correlation function has a feature that it becomes divergence-free, implying the lack of conformal invariance.
preprint: YITP-25-43, RIKEN-iTHEMS-Report-25
1 Introduction
@@
The origin of magnetism has been one of the greatest mysteries in nature since the days of ancient Greece and China. Even after the development of Maxwell’s theory of electromagnetism and classical statistical mechanics, ferromagnetism remains a mystery. As Bohr argued in 1911 in his PhD thesis, since a static magnetic field does not work on charged particles there can be no net magnetism in statistical equilibrium. It was then Heisenberg who, in 1928, brought the brilliant idea that the origin of magnetism must be quantum mechanical Heisenberg:1928mqa : It is not the electromagnetic interaction but the quantum exchange effect that will cause the alignment of spins and explain ferromagnetism. His model, the Heisenberg model, has been the starting point for understanding ferromagnetism in nature from statistical mechanics. It provides the simplest statistical models to study the universal nature of Currie’s phase transition with the help of the renormalization group, Monte Carlo simulation, and conformal bootstrap. This is a familiar story.
@@
In 1973, Aharony and Fisher challenged the above conventional wisdom Aharony_1 ; Aharony_2 . They pointed out that the long-range electromagnetic dipolar exchange between magnetization vectors, rather than the short-range quantum mechanical exchange effect of Heisernerbg, should lead to a relevant effect in the renormalization group sense although it is typically too tiny to explain the ferromagnetism itself. Accordingly, they further argue that it can drastically change the nature of Currie’s phase transition because in the theory of renormalization group, it is not the smallness of the bare parameter but whether it is relevant or irrelevant will determine the fixed point and nature of the phase transition. The salient features of the new renormalization group fixed point caused by the long-range electromagnetic dipolar exchange effect, which we call dipolar fixed point, are (1) it demands the magnetization vector fields to be constrained to be transverse, (2) as a consequence it is scale-invariant but not conformal invariant, (3) the critical exponents must be different from those of the Heisenberg fixed point.
@@
There are some experimental verifications of (1) and (2). Certain ferromagnetic materials such as Europium compounds EuO and EuS experimentally show the spin correlation functions with suppressed transverse components J.Kotzler_1986 , which is not only a clear signal of (1) but also that there is no conformal invariance as argued in Ref. Gimenez-Grau:2023lpz . Regarding (3), the measured critical exponents of EuO and EuS are close to those of the Heisenberg model PhysRevB.14.4908 , and the difference, if any, is yet to be seen.
@@
We here point out that there have not been as many theoretical works to determine the critical exponents of the dipolar fixed points as those of the Heisenberg fixed points. In addition to the one or two-loop results in the original paper by Aharony and Fisher, we have only three-loop computations in Ref. 2022NuPhB.98515990K and the functional renormalization group analysis in Ref. Nakayama:2023wrx . The situation should be in stark contrast with the case of the Heisenberg model, where the competition among the renormalization group Guida:1998bx ; 2001JMP….42…52J , Monte Carlo simulations Hasenbusch:2011zwv ; Hasenbusch:2020pwj , and conformal bootstrap Kos:2015mba ; Kos:2016ysd ; Chester:2020iyt are very severe, now culminating up to more than six digits. We should emphasize here that the dipolar fixed point is not conformal invariant and the conformal bootstrap is not applicable.
@@
Our main finding is that the measured critical exponents of the dipolar fixed points can be very different from those of the Heisenberg fixed points: $\nu=0.601(2)(^{+0}_{-2})$ and $\eta=0.132(8)(^{+0}_{-4})$ at $\lambda=8$, which should be compared with $\nu=0.712(1)(^{+3}_{-0})$ and $\eta=0.0318(27)(^{+10}_{-7})$ at $\lambda=0$. In Figure 1, we compare our results of $\nu$ with the previous Monte Carlo studies Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz as well as the conformal bootstrap Chester:2020iyt , the $\varepsilon$-expansion 2022NuPhB.98515990K , and the functional renormalization group Nakayama:2023wrx . As for $\lambda=0$, our result is almost consistent with the other numerical works, while for $\lambda=8$ it shows a clear difference from the one at $\lambda=0$. Thus, it strongly suggests that the dipolar fixed point realized at $\lambda=\infty$ would be in a different universality class from the Heisenberg fixed point.
Figure 1: Comparison plot on the critical exponent $\nu$ (for $\lambda=0$ and $8$) obtained in this paper and the previous studies by the Monte Carlo Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz (colored open symbols), the conformal bootstrap Chester:2020iyt (star-yellow symbol), the $\varepsilon$-expansion (filled symbols), and the functional renormalization group Nakayama:2023wrx (cross-blue-gray symbols). The filled circle-black symbols depict the 1-loop results of the $\varepsilon$-expansion obtained for $\lambda=0$ and $\infty$ in Ref. Aharony_1 and for $\lambda=8$ in this paper whereas the filled square-circle ones are the 3-loop results taken from Ref. 2022NuPhB.98515990K . The two results of the functional renormalization group are obtained with the LPA’ approximation, truncating the potential at the order $n=6$ or $16$ Nakayama:2023wrx .
@@
The rest of the paper is organized as follows. In Section 2, we present a field theoretical description of the Heisenberg model with the dipolar constraint and then discuss the predicted correlation function and the loop calculation of critical exponents. Furthermore, we also present a local description of the model, which gives a rather suitable formulation for lattice calculations. Then, we define the lattice Hamiltonian and observables for the simulation in Section 3 and explain the details of the simulation in Section 4. The effectiveness of the dipolar-constraint term is also discussed. In Section 5, we first explain the basic idea of our finite-scaling analysis and then show the main results of the critical temperature and exponents. The characteristic behavior of correlation functions near the critical temperature is discussed in Section 6. Section 7 is devoted to the summary and discussion. In Appendix A, we present the result of the renormalization group analysis. In Appendix B, we discuss the Hubbard–Stratonovich transformation of the model. In Appendix C, the fitting results of the magnetic susceptibility for determining its peak temperature are shown. The fitting results for interpolating $\beta$-dependent observables are shown in Appendix D. We discuss the system size and temperature dependences of the correlation function in Appendix E and F, respectively.
2 Heisenberg-dipolar model
2.1 Hamiltonian in the continuum theory
@@
From the dimensional analysis, this interaction is relevant. In the IR limit, we should arrive at a new fixed point called the dipolar fixed point. The fixed point has the peculiar property that it is scale-invariant but not conformal invariant Gimenez-Grau:2023lpz .
@@
The perturbative computation suggests that the value of the anomalous dimension is small; $\gamma_{\phi}\approx 0.0115\varepsilon^{2}$ Aharony_2 with $\Delta_{\phi}=\frac{d-2}{2}+\gamma_{\phi}$ in $4-\varepsilon$ dimensions, and the perturbative fixed point cannot be conformal invariant. We attempt the non-perturbative computation of $\gamma_{\phi}$ from the lattice Monte Carlo simulation in this work.
@@
Although it is not directly relevant to our work, we note that there is a hidden shift symmetry in this model. To see this let us use the Lagrange multiplier formalism to impose the constraint (5) within the local Hamiltonian. With the Lagrange multiplier field $U$, we add the constraint term $H_{\text{contstraint}}=\int d^{3}x\,U\partial_{i}\phi_{i}$. Physically, we may interpret $U$ as a magnetic potential. This added Hamiltonian has a shift symmetry, $U\to U+\text{const.}$., and we can use it to prove certain non-renormalization properties of the scale-invariant but non-conformal field theory Gimenez-Grau:2023lpz . Introducing the Lagrange multiplier on the Monte Carlo simulation, however, is difficult, so in the next subsection we further add the $U^{2}$ term to make the constraint into a (positive) cost term in the Hamiltonian at the expense of losing the shift symmetry.
2.2 The local description
@@
In the $\lambda\to 0$ limit, the result, $y_{t}=2-\varepsilon/2$, reproduces the one of the $O(4)$-vector model. On the other hand, in the $\lambda\to\infty$ limit, it agrees with the Aharony-Fisher result, $y_{t}=2-9\varepsilon/17$ Aharony_1 . Beyond two loops, $\lambda$ should be either 00 or $\infty$ at the fixed point but this one-loop estimate with variable $\lambda$ will be useful to discuss the comparison with the Monte Carlo simulations where $\lambda$ takes a finite value.
@@ 3 Lattice formula
This Hamiltonian is the Heisenberg model, adding the local cost function for the dipolar constraint  [^2] Our normalization of $\lambda$ when $J=-1$ here is chosen such that it directly corresponds to the continuum field-theory description in Section 2 within the mean-field theory analysis based on the Hubbard-Stratonovich transformation. See Appendix B..
@@
$M_{1},M_{2},M_{3}=\begin{pmatrix}J-2\lambda&0&0\\ -2\lambda&J&0\\ -2\lambda&0&J\end{pmatrix},\begin{pmatrix}J&-2\lambda&0\\ 0&J-2\lambda&0\\ 0&-2\lambda&J\end{pmatrix},\begin{pmatrix}J&0&-2\lambda\\ 0&J&-2\lambda\\ 0&0&J-2\lambda\end{pmatrix};$ (17)
@@
The magnetic susceptibility $\chi_{m}$ and the (4th order) Binder parameter $U$ are given by Binder:1981sa
@@
4 Simulation details
From now on, we fix $J=-1$ in the lattice Hamiltonian (14) and consider the cubic lattice with $L$ and and $V=L^{3}$ under the periodic boundary condition. The simulations are performed by changing $L$ from $6$ to $40$ in increments of $2$. We take mainly $\lambda=0$, namely the original Heisenberg model, and $\lambda=8$ for the local Heisenberg-dipolar model, which is shown to be regarded near the dipolar fixed point in Section 4.1.
@@ 4 Simulation details
4.1 Effectiveness of the local dipolar constraint
@@
5 Determination of critical exponents
5.1 Finite-size scaling and calculation strategy
To evaluate critical exponents from Monte Carlo simulations, there are several different approaches: (1) to study power laws of susceptibility $\chi_{m}(\beta)$ and specific heat $C(\beta)$ as a function of $(\beta-\beta_{c})$, (2) to study finite-size scaling of $U(\beta_{c})$ and $\chi_{m}(\beta_{c})$ at the critical temperature (3) to study correlation functions. Empirically, it is known that (2) gives the most precise (and hopefully most accurate) determination of critical exponents. For instance, in Ref. Hasenbusch:2020pwj , the author gives one of the state-of-art results on the critical exponents from a lattice model, which is the same universality-class model with the Heisenberg model. We essentially adopt this approach to obtain the exponents.
@@ 5.1 Finite-size scaling and calculation strategy
Due to the finite-size effects, $(\beta_{\times},U_{\times})$ depends on $L_{1}$ and $L_{2}$, but we can estimate its dependence from an appropriate finite-size correction ansatz. Here we consider a simple correction term of $O(L^{-\omega})$ with the exponent $\omega>0$ and assume the asymptotic behavior $U(\beta_{c},L)\simeq c_{\omega}L^{-\omega}+\mathrm{const.}$. Then, based on the scaling relation (24), we have an ansatz on the behavior of the Binder parameter around $\beta_{c}$ as in Ref. Binder:1981sa ,
@@
$f(L_{1},L_{2}):=-\frac{L_{2}^{-\omega}-L_{1}^{-\omega}}{L_{2}^{y_{t}}-L_{1}^{y_{t}}},\qquad g(L_{1},L_{2}):=\frac{L_{2}^{y_{t}}L_{1}^{-\omega}-L_{1}^{y_{t}}L_{2}^{-\omega}}{L_{2}^{y_{t}}-L_{1}^{y_{t}}},$ (27)
@@
Then, we obtain the exponent $\gamma$ using the relation $\gamma=\nu(2-\eta)$. The whole analysis of these procedures is performed using the jackknife method, which uses binned data and takes care of autocorrelation.
5.2 Intersecting points of the Binder parameter

### v3

# Monte Carlo study on Heisenberg model with local dipolar interaction

@@ Abstract
Aharony and Fisher showed that non-local dipolar effects in magnetism destabilize the Heisenberg fixed point in real ferromagnets, leading to a new fixed point, called the dipolar fixed point. The non-perturbative nature of the new fixed point, however, has not been uncovered for many decades. Inspired by the recent understanding that the dipolar fixed point is scale-invariant but not conformal invariant, we perform the Monte Carlo simulation of the local Heisenberg-dipolar model on the lattice of $40^{3}$ by introducing the local cost function parameterized by a parameter $\lambda$, and study its critical exponents, which should become identical to the dipolar fixed point of Aharony and Fisher in the infinite coupling limit $\lambda=\infty$. We find that the critical exponents become noticeably different from those of the Heisenberg fixed point for a finite coupling constant $\lambda=8$ (e.g. $\nu=0.601(2)(^{+0}_{-2})(^{+5}_{-4})_{\beta_{c}}$ in the local Heisenberg-dipolar model while $\nu=0.712(1)(^{+3}_{-0})(^{+1}_{-1})_{\beta_{c}}$ in the Heisenberg model), and the spin correlation function has a feature that it becomes divergence-free, implying the lack of conformal invariance.
preprint: YITP-25-43, RIKEN-iTHEMS-Report-25preprint: YITP-25-43, RIKEN-iTHEMS-Report-25
I Introduction
@@
The origin of magnetism has been one of the greatest mysteries in nature since the days of ancient Greece and China. Even after the development of Maxwell’s theory of electromagnetism and classical statistical mechanics, ferromagnetism remains a mystery. As Bohr argued in 1911 in his PhD thesis, since a static magnetic field does not work on charged particles there can be no net magnetism in statistical equilibrium. It was then Heisenberg who, in 1928, brought the brilliant idea that the origin of magnetism must be quantum mechanical [1]: It is not the electromagnetic interaction but the quantum exchange effect that will cause the alignment of spins and explain ferromagnetism. His model, the Heisenberg model, has been the starting point for understanding ferromagnetism in nature from statistical mechanics. It provides the simplest statistical models to study the universal nature of Currie’s phase transition with the help of the renormalization group, Monte Carlo simulation, and conformal bootstrap. This is a familiar story.
@@
In 1973, Aharony and Fisher challenged the above conventional wisdom [2, 3]. They pointed out that the long-range electromagnetic dipolar exchange between magnetization vectors, rather than the short-range quantum mechanical exchange effect of Heisernerbg, should lead to a relevant effect in the renormalization group sense although it is typically too tiny to explain the ferromagnetism itself. Accordingly, they further argue that it can drastically change the nature of Currie’s phase transition because in the theory of renormalization group, it is not the smallness of the bare parameter but whether it is relevant or irrelevant will determine the fixed point and nature of the phase transition. The salient features of the new renormalization group fixed point caused by the long-range electromagnetic dipolar exchange effect, which we call dipolar fixed point, are (1) it demands the magnetization vector fields to be constrained to be transverse, (2) as a consequence it is scale-invariant but not conformal invariant, (3) the critical exponents must be different from those of the Heisenberg fixed point.
@@
There are some experimental verifications of (1) and (2). Certain ferromagnetic materials such as Europium compounds EuO and EuS experimentally show the spin correlation functions with suppressed transverse components [4], which is not only a clear signal of (1) but also that there is no conformal invariance as argued in Ref. [5]. Regarding (3), the measured critical exponents of EuO and EuS are close to those of the Heisenberg model [6], and the difference, if any, is yet to be seen.
@@
We here point out that there have not been as many theoretical works to determine the critical exponents of the dipolar fixed points as those of the Heisenberg fixed points. In addition to the one or two-loop results in the original paper by Aharony and Fisher, we have only three-loop computations in Ref. [7] and the functional renormalization group analysis in Ref. [8]. The situation should be in stark contrast with the case of the Heisenberg model, where the competition among the renormalization group [9, 10], Monte Carlo simulations [11, 12], and conformal bootstrap [13, 14, 15] are very severe, now culminating up to more than six digits. We should emphasize here that the dipolar fixed point is not conformal invariant and the conformal bootstrap is not applicable.
@@
Our main finding is that the measured critical exponents of the dipolar fixed points can be very different from those of the Heisenberg fixed points: $\nu=0.601(2)(^{+0}_{-2})(^{+5}_{-4})_{\beta_{c}}$ and $\eta=0.132(8)(^{+0}_{-4})(^{+12}_{-11})_{\beta_{c}}$ at $\lambda=8$, which should be compared with $\nu=0.712(1)(^{+3}_{-0})(^{+1}_{-1})_{\beta_{c}}$ and $\eta=0.0318(27)(^{+10}_{-7})(^{+27}_{-28})_{\beta_{c}}$ at $\lambda=0$. In Figure 1, we compare our results of $\nu$ with the previous Monte Carlo studies [12, 11, 16, 17, 18, 19, 20, 21] as well as the conformal bootstrap [15], the $\varepsilon$-expansion [7], and the functional renormalization group [8]. As for $\lambda=0$, our result is almost consistent with the other numerical works, while for $\lambda=8$ it shows a clear difference from the one at $\lambda=0$. Thus, it strongly suggests that the dipolar fixed point realized at $\lambda=\infty$ would be in a different universality class from the Heisenberg fixed point.
Figure 1: Comparison plot on the critical exponent $\nu$ (for $\lambda=0$ and $8$) obtained in this paper and the previous studies by the Monte Carlo [12, 11, 16, 17, 18, 19, 20, 21] (colored open symbols), the conformal bootstrap [15] (star-yellow symbol), the $\varepsilon$-expansion (filled symbols), and the functional renormalization group [8] (cross-blue-gray symbols). The filled circle-black symbols depict the 1-loop results of the $\varepsilon$-expansion obtained for $\lambda=0$ and $\infty$ in Ref. [2] and for $\lambda=8$ in this paper whereas the filled square-circle ones are the 3-loop results taken from Ref. [7]. The two results of the functional renormalization group are obtained with the LPA’ approximation, truncating the potential at the order $n=6$ or $16$ [8].
@@
The rest of the paper is organized as follows. In Section II, we present a field theoretical description of the Heisenberg model with the dipolar constraint and then discuss the predicted correlation function and the loop calculation of critical exponents. Furthermore, we also present a local description of the model, which gives a rather suitable formulation for lattice calculations. Then, we define the lattice Hamiltonian and observables for the simulation in Section III and explain the details of the simulation in Section IV. The effectiveness of the dipolar-constraint term is also discussed. In Section V, we first explain the basic idea of our finite-scaling analysis and then show the main results of the critical temperature and exponents. The characteristic behavior of correlation functions near the critical temperature is discussed in Section VI. Section VII is devoted to the summary and discussion. In Appendix A, we present the result of the renormalization group analysis. In Appendix B, we discuss the Hubbard–Stratonovich transformation of the model. In Appendix C, the fitting results of the magnetic susceptibility for determining its peak temperature are shown. The fitting results for interpolating $\beta$-dependent observables are shown in Appendix D. We discuss the system size and temperature dependences of the correlation function in Appendix F and G, respectively.
II Heisenberg-dipolar model
II.1 Hamiltonian in the continuum theory
@@
From the dimensional analysis, this interaction is relevant. In the IR limit, we should arrive at a new fixed point called the dipolar fixed point. The fixed point has the peculiar property that it is scale-invariant but not conformal invariant [5].
@@
The perturbative computation suggests that the value of the anomalous dimension is small; $\gamma_{\phi}\approx 0.0115\varepsilon^{2}$ [3] with $\Delta_{\phi}=\frac{d-2}{2}+\gamma_{\phi}$ in $4-\varepsilon$ dimensions, and the perturbative fixed point cannot be conformal invariant. We attempt the non-perturbative computation of $\gamma_{\phi}$ from the lattice Monte Carlo simulation in this work.
@@
Although it is not directly relevant to our work, we note that there is a hidden shift symmetry in this model. To see this let us use the Lagrange multiplier formalism to impose the constraint (5) within the local Hamiltonian. With the Lagrange multiplier field $U$, we add the constraint term $H_{\text{contstraint}}=\int d^{3}x\,U\partial_{i}\phi_{i}$. Physically, we may interpret $U$ as a magnetic potential. This added Hamiltonian has a shift symmetry, $U\to U+\text{const.}$, and we can use it to prove certain non-renormalization properties of the scale-invariant but non-conformal field theory [5]. Introducing the Lagrange multiplier on the Monte Carlo simulation, however, is difficult, so in the next subsection we further add the $U^{2}$ term to make the constraint into a (positive) cost term in the Hamiltonian at the expense of losing the shift symmetry.
II.2 The local description
@@
In the $\lambda\to 0$ limit, the result, $y_{t}=2-\varepsilon/2$, reproduces the one of the $O(4)$-vector model. On the other hand, in the $\lambda\to\infty$ limit, it agrees with the Aharony-Fisher result, $y_{t}=2-9\varepsilon/17$ [2]. Beyond two loops, $\lambda$ should be either 0 or $\infty$ at the fixed point but this one-loop estimate with variable $\lambda$ will be useful to discuss the comparison with the Monte Carlo simulations where $\lambda$ takes a finite value.
@@ III Lattice formula
This Hamiltonian is the Heisenberg model, adding the local cost function for the dipolar constraint  [^2] Our normalization of $\lambda$ when J=1J=-1 here is chosen such that it directly corresponds to the continuum field-theory description in Section II within the mean-field theory analysis based on the Hubbard-Stratonovich transformation. See Appendix B..
@@
$M_{1},M_{2},M_{3}$
$=\begin{pmatrix}J-2\lambda&0&0\\ -2\lambda&J&0\\ -2\lambda&0&J\end{pmatrix},\begin{pmatrix}J&-2\lambda&0\\ 0&J-2\lambda&0\\ 0&-2\lambda&J\end{pmatrix},\begin{pmatrix}J&0&-2\lambda\\ 0&J&-2\lambda\\ 0&0&J-2\lambda\end{pmatrix}$; (17)
@@
The magnetic susceptibility $\chi_{m}$ and the (4th order) Binder parameter $U$ are given by [24]
@@
IV Simulation details
From now on, we fix J=1J=-1 in the lattice Hamiltonian (14) and consider the cubic lattice with $L$ and and $V=L^{3}$ under the periodic boundary condition. The simulations are performed by changing $L$ from $6$ to $40$ in increments of $2$. We take mainly $\lambda=0$, namely the original Heisenberg model, and $\lambda=8$ for the local Heisenberg-dipolar model, which is shown to be regarded near the dipolar fixed point in Section IV.1.
@@ IV Simulation details
IV.1 Effectiveness of the local dipolar constraint
@@
V Determination of critical exponents
V.1 Finite-size scaling and calculation strategy
To evaluate critical exponents from Monte Carlo simulations, there are several different approaches: (1) to study power laws of susceptibility $\chi_{m}(\beta)$ and specific heat $C(\beta)$ as a function of $(\beta-\beta_{c})$, (2) to study finite-size scaling of $U(\beta_{c})$ and $\chi_{m}(\beta_{c})$ at the critical temperature (3) to study correlation functions. Empirically, it is known that (2) gives the most precise (and hopefully most accurate) determination of critical exponents. For instance, in Ref. [12], the author gives one of the state-of-art results on the critical exponents from a lattice model, which is the same universality-class model with the Heisenberg model. We essentially adopt this approach to obtain the exponents.
@@ V.1 Finite-size scaling and calculation strategy
Due to the finite-size effects, $(\beta_{\times},U_{\times})$ depends on $L_{1}$ and $L_{2}$, but we can estimate its dependence from an appropriate finite-size correction ansatz. Here we consider a simple correction term of $O(L^{-\omega})$ with the exponent $\omega>0$ and assume the asymptotic behavior $U(\beta_{c},L)\simeq c_{\omega}L^{-\omega}+\mathrm{const.}$. Then, based on the scaling relation (24), we have an ansatz on the behavior of the Binder parameter around $\beta_{c}$ as in Ref. [24],
@@
$f(L_{1},L_{2})$ $=-\frac{L_{2}^{-\omega}-L_{1}^{-\omega}}{L_{2}^{y_{t}}-L_{1}^{y_{t}}},$ (27)
$g(L_{1},L_{2})$ $=\frac{L_{2}^{y_{t}}L_{1}^{-\omega}-L_{1}^{y_{t}}L_{2}^{-\omega}}{L_{2}^{y_{t}}-L_{1}^{y_{t}}},$
@@
Then, we obtain the exponent $\gamma$ using the relation $\gamma=\nu(2-\eta)$.
Taking into account the autocorrelation time mentioned in Section IV, we measured observables with 100 and 1000 sweeps intervals for $\lambda=0$ and $8$, respectively. In other words, each measurement data can be regarded as almost independent. We perform more than $10^{5}$ measurements for each parameter set of ($\beta,L,\lambda$). To estimate statistical errors, the entire analysis of the procedures listed above is performed using the jackknife method with the binned data, where the measurement data for each simulation parameter is divided into 100 bins. In this way, autocorrelation and error propagation were fully taken into account in the evaluation of statistical errors.
V.2 Intersecting points of the Binder parameter